# Maxwell’s Demon at Work: Efficient Pruning by Leveraging Saturation of Neurons

### v1

@@ 3 Demon Pruning
Leveraging insights from Sec. 2, we introduce Demon Pruning (DemP), summarized in Algorithm 1. It relies on two key elements: regularization and noise, influencing the learning dynamics to yield sparse solutions. Higher sparsity levels are achieved solely by increasing the peak regularization parameter ($\lambda$), as highlighted in Fig. 2.
Specifically, DemP updates parameters in the form:
$w_{t+1}=w_{t}-\eta\left(\nabla L(w_{t})+\nabla R(\gamma,\lambda)\right)+\xi_{\mathrm{asym}}$ (1)
where $\nabla L(w_{t})$ denotes the gradient estimator of some base optimizer, and $R(\gamma,\lambda)$ and $\xi_{\mathrm{asym}}$ are regularization and noise terms defined below.
@@
  Input: Learning rate $\eta$, pruning frequency $\tau$, regularization parameter $\lambda$, noise variance $\sigma^{2}$, subset of parameters with added scheduled regularization $\gamma$
  Initialize: Weights $w_{0}$, schedules $S_{\lambda}(t)$ and $S_{\sigma}(t)$
@@
     Update $\lambda_{t}$ and $\sigma_{t}^{2}$:$\lambda_{t}\leftarrow S_{\lambda}(t)$$\sigma_{t}^{2}\leftarrow S_{\sigma}(t)$
     Regularization term: $\nabla R(\gamma,\lambda_{t})$
@@
     Update weights:$w_{t+1}\leftarrow w_{t}-\eta(\nabla L(w_{t})+\nabla R(\gamma,\lambda_{t}))+\xi_{\mathrm{asym}}$
@@
        Remove dead neurons
@@
Regularization ($\bm{R(\gamma,\lambda})$):
Looking at the pre-activations ($\hat{\mathbf{z}}^{[l]}$) after batch normalization of the output of layer $l$ ($\mathbf{F}_{l}(\mathbf{W}^{[l]},{\bm{a}}^{[l-1]})$)
@@
hints toward two strategies to push $\hat{\mathbf{z}}^{[l]}$ toward zero: 1. regularizing all networks parameters, 2. regularizing the scale normalization parameters (${\bm{\gamma}}$). We refrain from regularizing the offset parameters (${\bm{\beta}}$) because negative offsets can help to drive the pre-activations into the negative region. Our method by default uses Lasso regularization on scale parameters of normalization layers, as in Liu et al. (2017a). Without normalization, weight regularization serves as an alternative to encourage sparsity (Fig. 2).
Based on the insights of the previous section, we use a one-cycled schedule (Smith & Topin, 2018) regularization strategy with minimal impacts:
  • Because neurons die early in training (see  2.1), regularization can be gradually decayed, allowing the network to recover some expressivity.
  • Previous works showcased that it can be beneficial to prune at an early stage instead of at initialization (Frankle et al., 2020; Rachwan et al., 2022). With our approach, this can be achieved by using warmup.
Employing a one-cycle strategy (Smith & Topin, 2018) effectively exploits the advantages of both warmup and decay phases. We illustrate the robustness of our design choices in the extensive ablation analysis in Sec. 4.3.
Asymmetric noise ($\bm{\xi_{\mathrm{asym}}}$):
Sec. 2.2 and Appendix B illustrate how SGD noise’s asymmetry aids in driving neurons towards inactivity by dissipating as they become inactive. DemP mimics this effect by injecting artificial noise exclusively into the weights of active units. Adding low variance Gaussian noise $\chi\sim\mathcal{N}(x|0,\sigma^{2})$ to the updates of active unit weights enhances sparsity without impacting performance (see Fig. 7). The relevance of injecting artificial noise is further justified in Fig. 13 where we show that it is by itself a sound strategy to sparsify neural networks, even in the absence of regularization.
DemP thus introduces two novel hyperparameters: growing regularization on normalization scaling and artificial noise. Notably, this addition does not affect standard hyperparameters, such as weight decay, maintaining DemP’s compatibility with existing training methodologies. By retaining this orthogonality, DemP seamlessly integrates into various training setups without any additional overhead. Furthermore, this approach allows for more extensive exploration of other hyperparameters to achieve sparser solutions.
Table 1: Comparison between different criteria when pruning a ResNet-50 trained on ImageNet with Adam around 80% (first line) and 90% (second line) weight sparsity. DemP improves test accuracy by 2.85% and 2.14% at those respective sparsities over the strongest baselines. Because structured pruning methods do not have precise control of weight sparsity, we report the closest numbers obtained to these target values. $\pm$ indicates the standard deviation, computed from 3 seeds. The sparsity numbers indicate the removed ratio.
Method Test accuracy Neuron sparsity Weight sparsity
Dense 74.98% ±$\pm 0.08$
Structured Methods SNAP 28.28% ±$\pm 0.08$ 36.9% 81.4%
27.17% ±$\pm 0.07$ 56.0% 90.1%
CroPit-S 28.34% ±$\pm 0.52$ 36.9% 81.4%
27.36% ±$\pm 0.16$ 53.2% 89.9%
EarlySNAP 68.67% ±$\pm 0.15$ 51.70% 80.37%
63.80% ±$\pm 0.58$ 66.6% 90.06%
EarlyCroP-S 68.26% ±$\pm 0.31$ 51.60% 79.97%
64.20% ±$\pm 0.27$ 66.6% 90.37%
DemP-L2 71.52% ±$\pm 0.09$ 61.83% 80.13%
66.34% ±$\pm 0.16$ 74.1% 89.93%
Table 2: Speedups and computational gains for the ResNet-50 model trained on ImageNet with Adam around 80%/90% weight sparsity, computed from 3 seeds. While SNAP and CroPit-S achieve better speedup, they do so by dropping significantly the test accuracy below 30%. DemP achieves much better speedups compared with EarlySNAP and EarlyCroP-S which maintain accuracy above 60%
Method Training time Training FLOPs Inference FLOPs
Dense 1.0x 1.0x (3.15e18) 1.0x (8.2e9)
SNAP 0.51x/0.48x 0.32x/0.25x 0.32x/0.25x
CroPit-S 0.52x/0.47x 0.32x/0.27x 0.32x/0.27x
EarlySNAP 0.95x/0.75x 0.63x/0.46x 0.63x/0.45x
EarlyCroP-S 0.94x/0.82x 0.66x/0.51x 0.66x/0.50x
DemP-L2 0.81x/0.61x 0.57x/0.42x 0.49x/0.34x
@@ 4 Empirical Evaluation
Through extensive empirical evaluation across various benchmarks, we consistently observe that DemP achieves superior performance/sparsity tradeoffs, particularly at high sparsity levels and when combined with Adam.
4.1 Setup
We focus our experiments on computer vision tasks, which is standard in pruning literature (Gale et al., 2019). We train ResNet-18 and VGG-16 networks on CIFAR-10, and ResNet-50 networks on ImageNet (He et al., 2016; Simonyan & Zisserman, 2015; Krizhevsky et al., 2009; Deng et al., 2009). We follow the training regimes from Evci et al. (2020) for ResNet architectures and use a setting similar to Rachwan et al. (2022) for the VGG to broaden the scope of our experiments. The results for pruning the VGG are reported in Appendix G and training details are provided in Appendix H.
Our method is a structured pruning method, removing entire neurons at once, and pruning happens during training, transitioning from a dense network to a sparse one. The methods we compare with follow this paradigm, except for approaches like Lasby et al. (2023), which achieves remarkable performance but primarily relies on unstructured pruning followed by structured reorganization. We employ the following structured pruning baselines: Crop-it/EarlyCrop (Rachwan et al., 2022), SNAP (Verdenius et al., 2020) and a modified version using the early pruning strategy from Rachwan et al. (2022) (identified as EarlySNAP). We trained these baselines using the recommended configuration of the original authors; in particular, we did not subject them to the regularization schedule employed by our method.
@@ 4.2 Results
DemP demonstrated superior performance when paired with Adam optimizer, consistently outperforming all baselines at high sparsity across the board. The margin in test error can reach up to 2.25% when training on CIFAR-10 and ImageNet (Fig. 4, 15 and Table 1). The results are more contrasted with SGDM: DemP performs similarly to baselines on ResNet-18 benchmarks (see Fig. 4, 15), is outperformed at higher sparsities on VGG (see Fig. 14) but slightly outperforms baselines when training a ResNet-50 on ImageNet (refer to Table 3). Additionally, DemP provides significant speedup when training on Imagenet with both optimizers, surpassing baselines yet again (Table 2 and  3).
Furthermore, as shown in Fig. 15, these conclusions hold when a ResNet-18 is trained with Leaky ReLU activations instead of ReLU, highlighting DemP’s flexibility across different activations functions.
Figure 5: We studied the impact of different schedules over the regularization parameter for our method, settling down on a one-cycle scheduler as default. Experiments were performed with ResNet-18 on CIFAR-10, with Lasso($\gamma$) regularization, across 3 seeds. Higher sparsities are obtained by increasing the peak strength of the added scheduled regularization. Left: With Adam optimizer. Right: With SGDM optimizer.
Figure 6: ResNet-18 networks trained on CIFAR-10 with different added regularization strategies, over 3 seeds. $\cdot(\gamma)$ denotes when regularization is only applied on the scale parameters of the normalization layer. Higher sparsities are obtained by increasing the peak strength of the added scheduled regularization. Left: With Adam, using L2($\gamma$) regularization slightly outperforms other strategies. Right: Using SGDM, the differences in performance become more pronounced, with Lasso regularization applied to scale parameters offering the optimal balance between sparsity and performance.
Figure 7: Left: ResNet-18 networks trained on CIFAR-10 with SGDM over 3 seeds. Adding a very small amount of Gaussian noise ($\sigma^{2}=5\times 10^{-5})$ to the live neurons provides the best tradeoff—performance is improved or maintained across all sparsity ratios. Right: The addition of noise proves very impactful in increasing sparsity without affecting performance on a ResNet-50 train on ImageNet. Each symbol represents a distinct peak regularization value for scale parameters, with one data point corresponding to each seed.
Figure 8: Measuring final accuracy vs. sparsity when dynamic pruning is enabled or not. Different symbols were associated with different regularization strengths. Experiments were performed with ResNet-18 on CIFAR-10 and Lasso($\gamma$) regularization for 3 seeds. Left: With Adam, there are barely any variations between the two strategies, allowing us to conclude that using dynamic pruning does not affect performance. Right: With SGDM, there is a gap in sparsity, but very little variation in performance. This also makes dynamic pruning an effective tool to achieve better sparsity tradeoffs with SGDM, in addition to training speedup.
@@ 4.3 Ablation
Regularizers. While the methodology works with both L2 and Lasso regularization, and either over the entire parameter space or specifically to the scale normalization parameters ($\gamma$), we empirically found that Lasso regularization of the scaling worked best for SGDM (Fig 6). For Adam, L2 regularization of the scaling slightly outperforms Lasso regularization of the scaling (Fig 6). As such, normalizing solely the scale parameters is beneficial to performance. For simplicity, we opted for Lasso regularization of the scale parameters in all subsequent experiments. Nevertheless, in the absence of normalization layers, traditional regularization methods remain viable for inducing sparsity in the model
Dynamic Pruning. To realize computational gain during training, we prune dynamically the NN at every $k$ steps (with a default of $k=1000$). Dead neurons end up removed almost as they appear, not giving them any chance to be revived. This strategy allows to speed up the training with no significant change in performance as shown in Fig. 8. We note that this smooth gradual pruning process is compatible with our approach in part because there is no added cost for computing the pruning criterion.
Dead Criterion Relaxation. The definition we choose for a dead neuron asks for it to be inactive to the entire dataset. In practice, we found that this criterion could be relaxed and defaulted to using 512 (2048 for ImageNet) examples from the training dataset to measure the death state (measuring across multiple minibatches when necessary). Fig. 12 shows that using this proxy for tracking dead units is sufficient. Pruning interventions become highly efficient, incurring only the computational cost of a few forward propagation batches.
@@
Added noise. Adding asymmetric noise is particularly beneficial to further increase sparsity (see Fig. 7). The noise is kept small, with a peak variance at $\sigma^{2}=5\times 10^{-5}$ and following the same schedule as the added regularization for maximal effect. We believe it acts as the small additional push helping neurons lingering close to the inactive region boundary to cross it.
@@
We conducted experiments to assess the necessity of using asymmetric noise in DemP (refer to Fig. 13). Surprisingly, we found that when all other aspects of DemP remained constant (including dynamic pruning), asymmetric noise was not essential. We yet decided to keep the added noise asymmetric due to significant differences observed in other settings (Fig. 10, middle) where the impact of symmetric noise on dead neuron accumulation was considerably less pronounced due to neuron revival.
Weight Decay. Our method defaults back to traditional regularization, with a term added directly to the loss, as opposed to the weight decay scheme proposed by Loshchilov & Hutter (2019). By doing so, the adaptive term in optimizers takes into account regularization, and neurons move more quickly toward the inactive region. From a pruning perspective, it achieves a higher sparsity than weight decay for the same regularization strength.
Activation function.  Dead neurons, and thus the pruning mechanism behind our method, are naturally defined with ReLU activation functions, in which neurons can completely deactivate. However, multiple activation functions, such as Leaky ReLU (Maas et al., 2013), also exhibit a “soft” saturated region. We postulate that neurons firing solely from the saturated region do not contribute much to the predictions and can be considered almost dead. We test this hypothesis by employing our method in a network with Leaky ReLU activations (Fig. 15), removing neurons with only negative activation across a large minibatch. Again, our method can outperform other structured methods when using Adam and offers a similar performance with SGDM.
5 Related Works
While our method monitors activations (Hu et al., 2016) and leverages increasing regularization (Wang et al., 2021) on scale parameters (Liu et al., 2017a), DemP introduces new techniques for pruning, beginning with the artificial asymmetric noise injected during optimization to further promote sparsity (Eq. 1.) Also, with DemP, sparsity naturally arises from the learning dynamics, removing the need for defining pruning criteria and the need for discrete pruning interventions impacting training. Non-contributing neurons can be promptly pruned as they become inactive, without compromising the model’s present or future performance (as illustrated in Fig. 8). Recent unstructured pruning approaches manually tailor sparsity distribution across layers (Mocanu et al., 2018; Evci et al., 2020), while structured pruning mostly rely on uniform layer-wise pruning (Rachwan et al., 2022). DemP alleviates this by letting the learning dynamics dictate the sparsity distribution across layers.
DemP eliminates the need for meticulous pruning timing (Wang et al., 2021; Rachwan et al., 2022) or the creation of a heuristic-based pruning schedule (Lee et al., 2023). Acknowledging that very early pruning can be detrimental (Frankle et al., 2020; Rachwan et al., 2022), DemP employs easy-to-tune regularization schedules to influence when units die. Iterative pruning (Verdenius et al., 2020) during a single intervention is also unnecessary, again because units die gradually during the early training phase with DemP.
Structured Pruning. Pruning is used to reduce the size and complexity of neural networks by removing redundant or less important elements, be they neurons or weights while maintaining their performance (LeCun et al., 1989). Recent advances such as those based on the Lottery Ticket Hypothesis (Frankle & Carbin, 2019) have demonstrated the existence of subnetworks trainable to comparable performance as their dense counterpart but with fewer parameters.
Demp is a structured pruning method that aims to remove entire structures within a network, such as channels, filters, or layers. It results in smaller and faster models that maintain compatibility with existing hardware accelerators and software libraries (Wen et al., 2016; Li et al., 2017). We highlight and benchmark against strong baselines that use criteria based on gradient flow to evaluate which nodes to prune (Verdenius et al., 2020; Wang et al., 2020; Rachwan et al., 2022). Other works employed either L0 or L1 regularization on gate parameters (or batch normalization scaling parameters) to enforce sparsity (Liu et al., 2017b; Louizos et al., 2018; You et al., 2019), but we do not benchmark them as they are outperformed by Rachwan et al. (2022).
Regularization-based pruning has been a popular approach, with canonical papers employing L0 or L1 regularization to induce sparsity directly (Louizos et al., 2018; Liu et al., 2017b; Ye et al., 2018) while L2 can help identify the connections to prune with the smallest weight criterion (Han et al., 2015). Because uniform regularization can quickly degrade performance (Wen et al., 2016; Lebedev & Lempitsky, 2016), Ding et al. (2018) and (Wang et al., 2019) proposed to adapt the regularization for different parameter groups. Recently, Wang et al. (2021) showed that growing the L2 regularization can leverage Hessian information to identify the filters to prune in pre-trained networks.
Dead Neurons. It is widely recognized that neurons, especially in ReLU networks, can saturate during training (Agarap, 2018; Trottier et al., 2017; Lu et al., 2019). In particular, Evci (2018) noted the connection of the dying rate with the learning rate and derived a pruning technique from it.
Mirzadeh et al. (2023), contemporary to our work, highlighted the strategic use of ReLU activations to achieve inference efficiency in transformer-based models. They propose an approach leveraging post-activation sparsity to improve inference speed in large language models.
Dead neurons are also studied in continual and reinforcement learning through the lens of plasticity loss (Berariu et al., 2021; Lyle et al., 2022), that progressively makes a model less capable of adapting to new tasks (Kirkpatrick et al., 2016). In some scenarios, a cause of plasticity loss has been attributed to the accumulation of dead units (Sokar et al., 2023; Lyle et al., 2023; Abbas et al., 2023; Dohare et al., 2021).

### v2

@@ 3 Demon Pruning
Leveraging insights from Section 2, we introduce Demon Pruning (DemP), a structured pruning method that dynamically induces neuron saturation during training to achieve high levels of sparsity. DemP achieves this through two primary mechanisms: (1) a dynamic regularization schedule applied to normalization scale parameters and (2) asymmetric noise injection targeting active neurons. The combination of these mechanisms ensures sparse networks without compromising performance. DemP is summarized in Algorithm 1.
@@
Input: Learning rate $\eta$, pruning frequency $\tau$, regularization one-cycled schedule $\lambda_{t}$, noise variance one-cycled schedule $\sigma^{2}_{t}$, learnable weights ww—including normalization scale parameters $\gamma$
Initialize: weights $w_{0}$—including normalization scale parameters $\gamma_{0}$
@@
  Regularization term: $\nabla R(\gamma_{t};\lambda_{t})$ affecting only the normalization scale parameters
@@
[…]
@@
   Prune dead neurons
@@
3.1 Regularization Schedule
[…]
@@ 3.1 Regularization Schedule
suggests two strategies to push $\hat{\mathbf{z}}^{[l]}$ toward zero: (1) regularizing all networks parameters, (2) regularizing the scale normalization parameters ($\bm{\gamma}$}). We refrain from regularizing the offset parameters ($\bm{\beta}$}), to allow for large negative offsets that can help to get negative pre-activations, leading to dead neurons post ReLU activation. By default, our method employs $\ell_{1}$-regularization $R({\bm{\gamma}};\lambda)=\lambda\|{\bm{\gamma}}\|_{1}$ on the scale parameters of normalization layers, as in liu2017learning, with a small modification. Based on the insights of Section 2, we adopt a one-cycled schedule (smith2018superconvergence) $\lambda_{t}$ for the regularization strength. This schedule involves an initial warmup phase with a linear growth to a peak value $\lambda$, followed by a cosine decay to fine-tune sparsity levels while maintaining performance.
Figure 3: Top: For ResNet-18 networks on CIFAR-10 trained with Adam (ReLU), DemP can find sparser solutions while maintaining better performance than other structured approaches. Higher levels of sparsity with DemP are obtained by increasing the peak strength of the added scheduled regularization. Bottom: With SGDM, DemP performance is comparable, without significant differences between methods. Left: Neural sparsity, structured methods. Right: Weight sparsity, structured methods.
Such a one-cycled strategy—originally applied to learning rates—effectively exploits the benefits of both warmup and decay phases (Fig. 4). While regularizing all network parameters is a viable alternative for inducing sparsity, applying regularization exclusively to normalization scale parameters proves more effective and robust, as shown in Fig. 2 and in Section 4.2.
3.2 Noise Injection
To promote neuron death, DemP injects Gaussian noise selectively into active weights during training. At each iteration, the noise is sampled as $\xi_{\mathrm{asym}}\sim\mathcal{N}(0,\sigma_{t}^{2})$ where the variance $\sigma_{t}^{2}$ follows the same one-cycled schedule as the regularization strength. The targeted noise injection enhances sparsity without degrading performance (see Fig. 6). The effectiveness of artificial noise is further validated in Fig. 13 where it demonstrates the ability to sparsify networks even in the absence of regularization.
3.3 Dynamic Pruning
DemP leverages targeted regularization and asymmetric noise to progressively accumulate inactive neurons, as described in Section 2. Inactive neurons are identified as those with outputs consistently equal to 0 across the dataset—a criterion relaxed in practice to minibatches (Fig. 12). For Leaky ReLU and GeLU networks, neurons with consistently negative activations are similarly deemed inactive.
Instead of deferring pruning until training concludes, DemP dynamically prunes dead neurons every $t$ iterations (default: 5000). This approach accelerates training by removing pruned neurons from the computation graph, reducing the embedding dimension. Importantly, it allows non-ReLU networks to recover without affecting final performance or sparsity levels (Fig. 7).
DemP thus introduces two new hyperparameters: scheduled regularization on normalization scale parameters and artificial noise. These additions are orthogonal to standard hyperparameters like weight decay, ensuring compatibility with existing training pipelines. By retaining this orthogonality, DemP integrates seamlessly into diverse setups, while leaving room for exploration of additional hyperparameters to achieve greater sparsity. Additional implementation details and ablations are provided in Appendix F.
@@ 4 Empirical Evaluation
Through extensive empirical evaluation across various benchmarks, we consistently observe that DemP achieves superior performance-sparsity tradeoffs, particularly at high sparsity levels and when combined with Adam. Comparing with SOTA DST methods, such as ChaseLFSHWMMPL23 and Lasby2023srigl, that perform restructuration exposes the benefits of DemP for achieving GPU-supported training FLOPs reduction (Table  2).
Setup: We focus our experiments on computer vision tasks, which is standard in pruning literature (Gale2019_state_of_sparsity). We train ResNet-18 and VGG-16 networks on CIFAR-10, and ResNet-50 networks on ImageNet (DBLP:conf/cvpr/HeZRS16; DBLP:journals/corr/SimonyanZ14a; krizhevsky2009learning; DBLP:conf/cvpr/DengDSLL009). We follow the training regimes from evci2020rigging for ResNet architectures and use a setting similar to DBLP:conf/icml/RachwanZCGAG22 for the VGG to broaden the scope of our experiments. The results for pruning the VGG are reported in Appendix G and training details are provided in Appendix I.
Table 1: Results when pruning a ResNet-50 trained on ImageNet with Adam at approximately 80% and 90% weight sparsity. DemP improves test accuracy by 2.85% and 2.14% at the respective sparsities. Because structured pruning methods lack precise control over weight sparsity levels., we report the closest numbers obtained to these target values. SNAP and CroPit-S achieve better speedup but do so by reducing the test accuracy below 30%. DemP achieves better speedups compared with methods maintaining accuracy above 60%. Standard deviation ($\pm$) was computed over three seeds.
[…]
Table 2: Results for pruning a ResNet-50 trained on ImageNet with SGDM at approximately 80% and 90% weight sparsity, formatted similarly to Table 1. DemP outperforms comparable structured pruning methods and is also compared with selected unstructured and restructured approaches, including recent DST methods like RigL, SRigL and Chase. While these methods achieve better sparsity-performance tradeoffs in terms of theoretical FLOP reduction, their sparsification is not fully GPU-amenable, leading to minimal training and inference speedup, as shown in the last column.
Method Test accuracy Neuron sparsity Weight sparsity Training time Training FLOPs Inference FLOPs GPU Inf. FLOPs
[…]
Dense 76.67% - - - - -
[…]
values obtained from jaxpruner
*values obtained from evci2020rigging
¶values obtained from Lasby2023srigl, neuron sparsity estimated from Fig. 3b
§values obtained from ChaseLFSHWMMPL23
Our method is a structured pruning approach that removes entire neurons during training, transitioning from a dense network to a sparse one. The methods we compare with follow this paradigm. We employ the following structured pruning baselines: Crop-it/EarlyCrop (DBLP:conf/icml/RachwanZCGAG22), SNAP (verdenius2020snipit) and a modified version using the early pruning strategy from DBLP:conf/icml/RachwanZCGAG22 (identified as EarlySNAP). We trained these baselines using the recommended configuration of the original authors; in particular, we did not apply the regularization schedule used in our method.
@@ 4.1 Results
[…]
Furthermore, as shown in Fig. 15, these conclusions hold when a ResNet-18 is trained with Leaky ReLU activations instead of ReLU, highlighting DemP’s adaptability across different activation functions.
Figure 4: Impact of different schedules over the regularization parameter for DemP, concluding that a one-cycle scheduler is a good default choice. Experiments were performed with ResNet-18 on CIFAR-10, with Lasso($\gamma$) regularization, across three seeds. Higher sparsities are obtained by increasing the peak strength of the added scheduled regularization. Left: With Adam optimizer. Right: With SGDM optimizer.
Figure 5: ResNet-18 networks trained on CIFAR-10 with different added regularization strategies, over three seeds. $\cdot(\gamma)$ denotes when regularization is only applied on the scale parameters of the normalization layer. Higher sparsities are obtained by increasing the peak strength of the added scheduled regularization. Left: With Adam, using L2($\gamma$) regularization slightly outperforms other strategies. Right: Using SGDM, the differences in performance become more pronounced, with Lasso regularization applied to scale parameters providing a favorable balance between sparsity and performance.
Figure 6: Left: ResNet-18 networks trained on CIFAR-10 with SGDM over three seeds. Adding a very small amount of Gaussian noise ($\sigma^{2}=5\times 10^{-5})$ to the live neurons offers the best balance between performance and sparsity. Right: The addition of noise proves very impactful in increasing sparsity without affecting performance on a ResNet-50 trained on ImageNet. Each symbol represents a distinct peak regularization value for scale parameters, where each data point represents a distinct seed.
@@ 4.2 Ablation
Regularizers. While the methodology works with both L2 and Lasso regularization, and either applied to the entire parameter space or specifically to the scale normalization parameters ($\gamma$), we empirically found that Lasso regularization of the scaling worked best for SGDM (Fig. 5). For Adam, L2 regularization of the scaling slightly outperforms Lasso regularization of the scaling (Fig. 5). As such, applying regularization solely to the scale parameters is beneficial to performance. For simplicity, we opted for Lasso regularization of the scale parameters in all subsequent experiments. Nevertheless, in the absence of normalization layers, traditional regularization methods remain viable for inducing sparsity in the model.
Dynamic Pruning. To realize computational gain during training, we dynamically prune the NN at every $t$ steps (with a default of $t=5000$). Dead neurons are removed almost as soon as they appear, preventing their revival. This strategy enables faster training with no significant change in performance as shown in Fig. 7. We note that this smooth gradual pruning process is compatible with our approach in part because there is no added cost for computing the pruning criterion.
@@
Figure 7: Measuring the impact of dynamic pruning on the accuracy-sparsity tradeoff. The symbols denote different regularization strengths. Experiments were performed with ResNet-18 on CIFAR-10 and Lasso($\gamma$) regularization (three seeds). Left: With Adam, there are minimal differences between the two strategies, indicating that dynamic pruning does not affect performance. Right: With SGDM, dynamic pruning further improves sparsity without degrading performance, while also providing training speedup.
@@
Added noise. Adding asymmetric noise is particularly beneficial to further increase sparsity (see Fig. 6). The noise is kept small, with a peak variance at $\sigma^{2}=5\times 10^{-5}$ and following the same schedule as the added regularization for maximal effect. We believe it acts as the small additional push helping neurons that linger close to the inactive region boundary to cross it.
@@
We conducted experiments to assess the necessity of using asymmetric noise in DemP (Fig. 13). We found that when all other aspects of DemP remained constant (including dynamic pruning), asymmetric noise was not strictly essential. Yet, we note that dynamic pruning mimics the effect of asymmetric noise: removal of dead neurons also removes the possibility of random revival. We decided to keep the added noise asymmetric due to significant differences observed in other settings (Fig. 9, middle) where symmetric noise reduces dead neuron accumulation considerably due to neuron revival.
Additional ablations on the death criterion relaxation and weight decay are discussed respectively in Appendix F.3 and Appendix F.5. We also discuss the choice of activation functions in Appendix F.7, showing that DemP offers similar performance when using Leaky ReLU activations (Fig. 15).
Table 3: Using DemP to prune the MLP layers of a ViT-B/16 model. Solely by injecting artificial noise ($\lambda=0$), DemP removes up to 90% of the neurons. We report the neural sparsity across the MLP layers but provide the equivalent weight sparsity for the entire network.
Method Test accuracy Neuron sparsity Weight sparsity Training time
[…]
4.3 Transformers
To assess the compatibility of DemP with transformer-based architectures, we conducted exploratory experiments on Vision Transformer (ViT) networks (VisionTransformersDosovitskiyB0WZ21). Our focus was on pruning the fully connected layers because they contain the majority of the parameters and use GELU activation functions. Notably, we observed a substantial number of neurons dying during regular training, which led us to explore DemP’s performance without regularization ($\lambda=0$), retaining only the added noise.
The exploratory results for a ViT-B/16 model trained on ImageNet are presented in Table 3. We found that 90% of the fully connected neurons were readily pruned using DemP. Additionally, replacing the GELU activations with ReLU (relustrikebackMirzadehAMMTSRF) resulted in only a slight performance degradation while still allowing the pruning of 43% of the neurons. Further details are provided in Appendix G.2.